# A Definition of AGI

### v2

@@ 1 Introduction
To systematically investigate whether AI systems possess this spectrum of abilities, we ground our approach in the Cattell-Horn-Carroll (CHC) theory of cognitive abilities (Carroll, 1993; McGrew, 2009; Schneider and McGrew, 2018; McGrew, 2023; McGrew et al., 2023), the most empirically validated model of human intelligence. CHC theory is primarily derived from the synthesis of over a century of iterative factor analysis of diverse collections of cognitive ability tests. In the late 1990’s to 2000’s almost all major clinical, individually administered tests of human intelligence have iterated towards test revisions that were either explicitly or implicitly based on CHC model test design blueprints (Keith and Reynolds, 2010; Schneider and McGrew, 2018). CHC theory provides a hierarchical taxonomic map of human cognition. It breaks down general intelligence into distinct broad abilities and numerous narrow abilities (such as induction, associative memory, or spatial scanning). Readers interested in the strengths and limitations of the CHC framework are directed to further scholarly discussions (Wasserman, 2019; Canivez and Youngstrom, 2019).
@@ 11 Auditory Processing (A)
The table summarizes current AI system performance on Auditory Processing (A) tasks. GPT-4 had no ability to audio processing capabilities, while GPT-5’s capabilities are appreciable but incomplete.
@@ 13 Discussion
Working Memory vs. Long-Term Storage: A prominent contortion is the reliance on massive context windows (Working Memory, WM) to compensate for the lack of Long-Term Memory Storage (MS). Practitioners use these long contexts to manage state and absorb information (e.g., entire codebases). However, this approach is inefficient, computationally expensive, and can overload the system’s attentional mechanisms. It ultimately fails to scale for tasks requiring days or weeks of accumulated context. A long-term memory system might take the form of a module (e.g., a LoRA adapter (Hu et al., 2021)) that continually adjusts model weights to incorporate experiences.
@@
The Engine Analogy. Our multifaceted view of intelligence suggests an analogy to a high-performance engine, where overall intelligence is the “horsepower” (Jensen, 2000). An artificial mind, much like an engine, is ultimately constrained by its weakest components. See Figure˜3 to understand the relations between these abilities. Currently, several critical parts of the AI “engine” are highly defective. This severely limits the overall “horsepower” of the system, regardless of how optimized other components might be. This framework identifies these defects to guide our assessment and how far we are from AGI.
@@
Solving the Dataset vs. Solving the Task. Our operationalization relies on task specifications. We occasionally elaborate on these task specifications with specific datasets, and we usually treat them as necessary but not sufficient for solving the task. Moreover, solving our illustrative examples do not imply the task is solved, as our collection of examples are not exhaustive. It is the default for automatic evaluations to inadequately cover their target phenomena (Yogatama et al., 2019), so our operationalization is far more likely to be robust and stand the test of time compared to existing automated evaluations. Since we couch our definition in a collection of tasks rather than heavily depend on specific existing datasets, we can test AI systems using the best available tests at the time.
@@
Related Work. Ilić and Gignac (2024) and Ren et al. (2024) find that a variety of AI systems’ capabilities are highly correlated with pre-training compute. Gignac and Szodorai (2024) discuss human psychometrics and testing the intelligence of AI systems. Turing (1950) argues that the Turing Test can indicate general ability. Gubrud (1997) proposed an early definition of AGI in 1997. Marcus et al. (2016) discuss the need to move beyond the Turing Test to capture the multidimensional nature of intelligence. Morris et al. (2023) articulate levels of AGI based on performance percentiles. Legg and Hutter (2007) discuss various tests for general machine intelligence.
Limitations. First, our conceptualization of intelligence is not exhaustive. It deliberately excludes certain faculties, such as the kinesthetic abilities proposed in alternative frameworks like Gardner’s theory of multiple intelligences (Gardner, 1993). Second, our illustrative examples are specific to the English language and are not culturally agnostic. Future research could involve adapting these tests across diverse linguistic and cultural contexts. Furthermore, our operationalization has inherent constraints. The General Knowledge (K) tests are necessarily selective and do not assess the full breadth of possible subject areas. A 100% AGI score represents a “highly proficient” well-educated individual who has achieved mastery across these tested dimensions, rather than well-educated in the sense of having a college degree. Moreover, while the scoring weights we employ are necessary for quantitative measurement, they represent one of many possible configurations. We give equal weight to each broad ability (10%) to prioritize breadth, but more discretionary weighting schemes could be reasonable. The results are contingent on these methodological choices, and future work could explore alternative collections of tasks and weighting schemes. Finally, while the aggregate AGI Score is provided for convenience, it could be misleading. A simple summation can obscure critical failures in bottleneck capabilities. For example, an AI system with a 90% AGI Score but 0% on Long-Term Memory Storage (MS) would be functionally impaired by a form of “amnesia,” severely limiting its capabilities despite a high overall score. Therefore, we recommend reporting the AI system’s cognitive profile and not just its AGI Score.
@@
Recursive AI is an AI that can independently conducting the entire AI R&D lifecycle, leading to the creation of markedly more advanced AI systems without human input.
@@
Superintelligence is an AI that greatly exceeds the cognitive performance of humans in virtually all domains of interest (Bostrom, 2014).
@@
Our AGI definition is about human-level AI, not economically-valuable AI, nor economy-level AI. OpenAI and Microsoft have reportedly considered AGI to be an AI that can generate $100 billion in profit (TechCrunch, 2024). We do not conflate AGI with economically valuable AI because narrow technologies, such as the iPhone, can generate billions in economic value, despite not being generally intelligent. Meanwhile, Replacement AI is about economy-level AI, and it includes physical tasks, unlike AGI.
Recursive AI removes the need for human researchers and “closes the loop” on AI R&D, enabling rapid, recursive capability gains (an “intelligence recursion” (Hendrycks et al., 2025)) without human scientific input and could potentially lead to Superintelligence.
@@ Acknowledgments
We would like to thank Arunim Agarwal, Oliver Zhang, Anders Edson, and Matthew Blyth for their helpful feedback.
@@ B.2 Reading Comprehension
Sentence Level: Reliably solving Winograd schemas is sufficient for the 1%. For example, >85% accuracy on WinoGrande [Sakaguchi et al., 2019] strongly suggests proficiency.
@@ B.3 Writing Ability
Test: If the model receives a 4 or greater out of 6 on GRE Analytical Writing prompts [Zhong et al., 2025], then that is sufficient for 3%, subject to memorization and robustness checks.
@@ B.4 English Usage Knowledge
Test: For sentence-level English usage knowledge, it is necessary that AI systems be able to achieve greater than a 60% correlation on the Corpus of Linguistic Acceptability [Warstadt et al., 2019].
@@ C.1 Arithmetic
Test: Greater than 95% on GSM8K [Cobbe et al., 2021] is sufficient for the 2%, subject to memorization and robustness checks.
@@ C.2 Algebra
“The first three terms of a geometric sequence are the integers a, 720, b, where a < 720 < b. What is the sum of the digits of the least possible value of b? Answer choices: (A) 9, (B) 12, (C) 16, (D) 18, (E) 21” Answer: E (21) [Mathematical Association of America, 2024]
“Integers a, b, and c satisfy ab + c = 100, bc + a = 87, and ca + b = 60. What is ab + bc + ca? Answer choices: (A), 212 (B), 247 (C), 258 (D), 276 (E) 284” Answer: D (276) [Mathematical Association of America, 2024]
Test: Greater than 90% on MATH [Hendrycks et al., 2021b] Algebra is sufficient for the 2%, subject to memorization and robustness checks.
@@ C.3 Geometry
Test: Greater than 95% on MATH [Hendrycks et al., 2021b] Geometry is sufficient for the 2%, subject to memorization and robustness checks.
@@ C.4 Probability
“Suppose an airline accepted 12 reservations for a commuter plane with 10 seats. They know that 7 reservations went to regular commuters who will show up for sure. The other 5 passengers will show up with a 50% chance, independently of each other. (a) Find the probability that the flight will be overbooked, i.e., more passengers will show up than seats are available. (b) Find the probability that there will be empty seats. (c) Let X be the number of passengers turned away. Find E(X).” Answer: (a) 0.1875, (b) 0.5, (c) 0.219 [Pitman, 1993]
“Suppose N dice are rolled, where . (a) Given that no two of the N dice show the same face, what is the probability that one of the dice shows a six? Give a formula in terms of N. (b) In (a) the number of dice N was fixed, but now repeat assuming instead that N is random, determined as the value of another die roll. Your answer now should be simply a number, not involving N.” Answer: (a) N/6, (b) 0.3604 [Pitman, 1993]
@@ C.5 Calculus
Answer: C [College Board, n.d.]
“A circular object is increasing in size in some unspecified manner, but it is known that when the radius is 6, the rate of change of the radius is 4. Find the rate of change of the area when the radius is 6.” Answer: [Spivak, ]
@@ D.1 Deduction
“David knows Mr. Zhang’s friend Jack, and Jack knows David’s friend Ms. Lin. Everyone of them who knows Jack has a master’s degree, and everyone of them who knows Ms. Lin is from Shanghai. Who is from Shanghai and has a master’s degree? (A) David. (B) Jack. (C) Mr. Zhang. (D) Ms. Lin.” Answer: A [Liu et al., 2020]
“Last night, Mark either went to play in the gym or visited his teacher Tony. If Mark drove last night, he didn’t go to play in the gym. Mark would go visit his teacher Tony only if he and his teacher had an appointment. In fact, Mark had no appointment with his teacher Tony in advance. Which is true based on the above statements? (A) Mark went to the gym with his teacher Tony last night. (B) Mark visited his teacher Tony last night. (C) Mark didn’t drive last night. (D) Mark didn’t go to the gym last night.” Answer: C [Liu et al., 2020]
Test: An accuracy level of 86% (human-level) on LogiQA 2.0 [Liu et al., 2023] is sufficient for the 2%, subject to memorization and robustness checks.
@@ D.2 Induction
For induction tests, we use Raven’s Progressive Matrices (RPMs) [John and Raven, 2003]. As mentioned above, we do not treat RPMs as a strong indicator of overall AI system intelligence, rather a direct measurement of abstract inductive reasoning abilities.
@@
Test: This is related to the ARC-AGI challenge [Chollet, 2019].
@@ D.3 Theory of Mind
The can of Pringles has moldy chips in it. Mary picks up the can in the supermarket and walks to the cashier. Is Mary likely to be aware that “The can of Pringles has moldy chips in it.”? Answer: No. [Kim et al., 2023]
@@ E.1.2 Transformation Sequence
Test: A related benchmark is Michelangelo [Vodrahalli et al., 2024], but it is substantially harder since it uses very long sequences, whereas we use shorter sequences.
@@ G.1.6 Figural Fluency
Draw as many unique designs as possible by connecting the dots with exactly four straight lines in 60 seconds [Delis et al., 2001]:
@@ G.2 Hallucinations
Test: The hallucination rate of the system on SimpleQA [Wei et al., 2024] (without tools) must be less than 5%.
@@ H.1 Perception
Note: Image anomaly detection is highly related to the “Odd-Item Out” and the “What’s Missing (WHM)” RIAS-2 subtests [Reynolds and Kamphaus, 2015].
@@
What is this? Answer: Airplane [Real et al., 2017]
What does this depict? Answer: Siberian Husky [Hendrycks et al., 2021a]
What does this distorted image depict? Answer: Zebras [He et al., 2021]
@@
[Karpathy and Fei-Fei, 2015]
Create a descriptive caption for this. Answer: A ferret rests its head on a black remote control. [Karpathy and Fei-Fei, 2015]
@@
U Intuitivephysics1.mp4 Is this animated scene physically plausible? Answer: No. [Bordes et al., 2025]
U Intuitivephysics2.mp4 Is this animated scene physically plausible? Answer: Yes. [Bordes et al., 2025]
U Running.mov Is this animated scene anomalous? Answer: Yes. [vid, 2024]
@@
For video anomaly detection, it is necessary to receive over 85% on IntPhysics 2 [Bordes et al., 2025], a measurement of intuitive physics understanding.

### v3

@@ 1 Introduction
To systematically investigate whether AI systems possess this spectrum of abilities, we ground our approach in the Cattell-Horn-Carroll (CHC) theory of cognitive abilities (chc_theory; mcgrew2009chc; schneider2018chc; mcgrew2023carroll; mcgrew2023psychometric), the most empirically validated model of human intelligence. CHC theory is primarily derived from the synthesis of over a century of iterative factor analysis of diverse collections of cognitive ability tests. In the late 1990’s to 2000’s almost all major clinical, individually administered tests of human intelligence have iterated towards test revisions that were either explicitly or implicitly based on CHC model test design blueprints (Keith2010CattellHornCarrollAA; schneider2018chc). CHC theory provides a hierarchical taxonomic map of human cognition. It breaks down general intelligence into distinct broad abilities and numerous narrow abilities (such as induction, associative memory, or spatial scanning). Readers interested in the strengths and limitations of the CHC framework are directed to further scholarly discussions (wasserman2019; canivez2019).
@@ 11 Auditory Processing (A)
The table summarizes current AI system performance on Auditory Processing (A) tasks. GPT-4 had no audio processing capabilities, while GPT-5’s capabilities are appreciable but incomplete.
@@ 13 Discussion
Working Memory vs. Long-Term Storage: A prominent contortion is the reliance on massive context windows (Working Memory, WM) to compensate for the lack of Long-Term Memory Storage (MS). Practitioners use these long contexts to manage state and absorb information (e.g., entire codebases). However, this approach is inefficient, computationally expensive, and can overload the system’s attentional mechanisms. It ultimately fails to scale for tasks requiring days or weeks of accumulated context. A long-term memory system might take the form of a module (e.g., a LoRA adapter (hu2021loralowrankadaptationlarge)) that continually adjusts model weights to incorporate experiences.
@@
The Engine Analogy. Our multifaceted view of intelligence suggests an analogy to a high-performance engine, where overall intelligence is the “horsepower” (Jensen2000_gFactor). An artificial mind, much like an engine, is ultimately constrained by its weakest components. See Figure˜3 to understand the relations between these abilities. Currently, several critical parts of the AI “engine” are highly defective. This severely limits the overall “horsepower” of the system, regardless of how optimized other components might be. This framework identifies these defects to guide our assessment and how far we are from AGI.
@@
Solving the Dataset vs. Solving the Task. Our operationalization relies on task specifications. We occasionally elaborate on these task specifications with specific datasets, and we usually treat them as necessary but not sufficient for solving the task. Moreover, solving our illustrative examples do not imply the task is solved, as our collection of examples are not exhaustive. It is the default for automatic evaluations to inadequately cover their target phenomena (Yogatama2019LearningAE), so our operationalization is far more likely to be robust and stand the test of time compared to existing automated evaluations. Since we couch our definition in a collection of tasks rather than heavily depend on specific existing datasets, we can test AI systems using the best available tests at the time.
@@
Related Work. Ili2024 and ren2024safetywashing find that a variety of AI systems’ capabilities are highly correlated with pre-training compute. Gignac2024DefiningIB discuss human psychometrics and testing the intelligence of AI systems. turing1950computing argues that the Turing Test can indicate general ability. Gubrud1997AGI proposed an early definition of AGI in 1997. Marcus2016BeyondTT discuss the need to move beyond the Turing Test to capture the multidimensional nature of intelligence. jin2025crl connect factor analysis of human cognition to AI capabilities. Morris2023LevelsOA articulate levels of AGI based on performance percentiles. Legg2007TestsOM discuss various tests for general machine intelligence.
Limitations. First, our conceptualization of intelligence is not exhaustive. It deliberately excludes certain faculties, such as the kinesthetic abilities proposed in alternative frameworks like Gardner’s theory of multiple intelligences (Gardner1993MultipleIT). Second, our illustrative examples are specific to the English language and are not culturally agnostic. Future research could involve adapting these tests across diverse linguistic and cultural contexts. Furthermore, our operationalization has inherent constraints. The General Knowledge (K) tests are necessarily selective and do not assess the full breadth of possible subject areas. A 100% AGI score represents a “highly proficient” well-educated individual who has achieved mastery across these tested dimensions, rather than well-educated in the sense of having a college degree. Moreover, while the scoring weights we employ are necessary for quantitative measurement, they represent one of many possible configurations. We give equal weight to each broad ability (10%) to prioritize breadth, but more discretionary weighting schemes could be reasonable. The results are contingent on these methodological choices, and future work could explore alternative collections of tasks and weighting schemes. Finally, while the aggregate AGI Score is provided for convenience, it could be misleading. A simple summation can obscure critical failures in bottleneck capabilities. For example, an AI system with a 90% AGI Score but 0% on Long-Term Memory Storage (MS) would be functionally impaired by a form of “amnesia,” severely limiting its capabilities despite a high overall score. Therefore, we recommend reporting the AI system’s cognitive profile and not just its AGI Score.
@@
Pandemic AI is an AI that can engineer and produce new, infectious, and virulent pathogens that could cause a pandemic (li2024wmdp; vct).
@@
Recursive AI is an AI that can independently conduct the entire AI R&D lifecycle, leading to the creation of markedly more advanced AI systems without human input.
@@
Superintelligence is an AI that greatly exceeds the cognitive performance of humans in virtually all domains of interest (Bostrom2014Superintelligence).
@@
Our AGI definition is about human-level AI, not economically-valuable AI, nor economy-level AI. OpenAI and Microsoft have reportedly considered AGI to be an AI that can generate $100 billion in profit (techcrunch2024microsoftagi). We do not conflate AGI with economically valuable AI because narrow technologies, such as the iPhone, can generate billions in economic value, despite not being generally intelligent. Meanwhile, Replacement AI is about economy-level AI, and it includes physical tasks, unlike AGI.
Recursive AI removes the need for human researchers and “closes the loop” on AI R&D, enabling rapid, recursive capability gains (an “intelligence recursion” (Hendrycks2025SuperintelligenceSE)) without human scientific input and could potentially lead to Superintelligence.
@@ Acknowledgments
We would like to thank Arunim Agarwal, Oliver Zhang, Anders Edson, John Guan, and Matthew Blyth for their helpful feedback.
@@ A.1 Commonsense
System performance on PIQA [piqa] must exceed 85% accuracy.
System performance on ETHICS Commonsense Morality [ethics] must exceed 80% accuracy.
@@ B.2 Reading Comprehension
Sentence Level: Reliably solving Winograd schemas is sufficient for the 1%. For example, >85% accuracy on WinoGrande [winogrande] strongly suggests proficiency.
Paragraph Level: Model accuracy on COQA [coqa] must exceed 85%, ReCoRD [zhang2018recordbridginggaphuman] accuracy must exceed 90%, and LAMBADA [lambada] accuracy must exceed 80% (zero shot).
Document Level: Model accuracy exceeding 55% on LongBench v2 [longbench] suggests proficiency. Since models must determine if a question is underdetermined, it should also have a hallucination rate of less than 1% on Vectara HHEM [vectara].
@@ B.3 Writing Ability
Test: If the model receives a 4 or greater out of 6 on GRE Analytical Writing prompts [zhong2024evaluating], then that is sufficient for 3%, subject to memorization and robustness checks.
@@ B.4 English Usage Knowledge
Test: For sentence-level English usage knowledge, it is necessary that AI systems be able to achieve greater than a 60% correlation on the Corpus of Linguistic Acceptability [cola].
@@ C.1 Arithmetic
Test: Greater than 95% on GSM8K [gsm8k] is sufficient for the 2%, subject to memorization and robustness checks.
@@ C.2 Algebra
“The first three terms of a geometric sequence are the integers a, 720, b, where a < 720 < b. What is the sum of the digits of the least possible value of b? Answer choices: (A) 9, (B) 12, (C) 16, (D) 18, (E) 21” Answer: E (21) [amc10a2024]
“Integers a, b, and c satisfy ab + c = 100, bc + a = 87, and ca + b = 60. What is ab + bc + ca? Answer choices: (A), 212 (B), 247 (C), 258 (D), 276 (E) 284” Answer: D (276) [amc10a2024]
Test: Greater than 90% on MATH [math_dataset] Algebra is sufficient for the 2%, subject to memorization and robustness checks.
@@ C.3 Geometry
Test: Greater than 95% on MATH [math_dataset] Geometry is sufficient for the 2%, subject to memorization and robustness checks.
@@ C.4 Probability
“Suppose an airline accepted 12 reservations for a commuter plane with 10 seats. They know that 7 reservations went to regular commuters who will show up for sure. The other 5 passengers will show up with a 50% chance, independently of each other. (a) Find the probability that the flight will be overbooked, i.e., more passengers will show up than seats are available. (b) Find the probability that there will be empty seats. (c) Let X be the number of passengers turned away. Find E(X).” Answer: (a) 0.1875, (b) 0.5, (c) 0.219 [pitman_probability]
“Suppose N dice are rolled, where . (a) Given that no two of the N dice show the same face, what is the probability that one of the dice shows a six? Give a formula in terms of N. (b) In (a) the number of dice N was fixed, but now repeat assuming instead that N is random, determined as the value of another die roll. Your answer now should be simply a number, not involving N.” Answer: (a) N/6, (b) 0.3604 [pitman_probability]
@@ C.5 Calculus
“ is?
1
3
9
nonexistent ”
Answer: B [ap_calculus]
“Which of the following limits is equal to ?
Answer: D [ap_calculus]
@@
Answer: C [ap_calculus]
“A circular object is increasing in size in some unspecified manner, but it is known that when the radius is 6, the rate of change of the radius is 4. Find the rate of change of the area when the radius is 6.” Answer: [spivak_calculus]
@@ D.1 Deduction
“David knows Mr. Zhang’s friend Jack, and Jack knows David’s friend Ms. Lin. Everyone of them who knows Jack has a master’s degree, and everyone of them who knows Ms. Lin is from Shanghai. Who is from Shanghai and has a master’s degree? (A) David. (B) Jack. (C) Mr. Zhang. (D) Ms. Lin.” Answer: A [logiqa]
“Last night, Mark either went to play in the gym or visited his teacher Tony. If Mark drove last night, he didn’t go to play in the gym. Mark would go visit his teacher Tony only if he and his teacher had an appointment. In fact, Mark had no appointment with his teacher Tony in advance. Which is true based on the above statements? (A) Mark went to the gym with his teacher Tony last night. (B) Mark visited his teacher Tony last night. (C) Mark didn’t drive last night. (D) Mark didn’t go to the gym last night.” Answer: C [logiqa]
Test: An accuracy level of 86% (human-level) on LogiQA 2.0 [Liu2023LogiQA2] is sufficient for the 2%, subject to memorization and robustness checks.
@@ D.2 Induction
For induction tests, we use Raven’s Progressive Matrices (RPMs) [raven1938]. As mentioned above, we do not treat RPMs as a strong indicator of overall AI system intelligence, rather a direct measurement of abstract inductive reasoning abilities.
@@
Test: This is related to the ARC-AGI challenge [chollet2019measureintelligence].
@@ D.3 Theory of Mind
The can of Pringles has moldy chips in it. Mary picks up the can in the supermarket and walks to the cashier. Is Mary likely to be aware that “The can of Pringles has moldy chips in it.”? Answer: No. [fantom]
@@
An accuracy level at or above 87.5% (human-level) on FANToM [fantom] is necessary for the 2%.
An accuracy level at or above 85.4% (human-level) on ToMBench [chen2024tombenchbenchmarkingtheorymind] is necessary for the 2%.
@@ D.4 Planning
Tests:
An accuracy of 90% or above on Natural Plan [natural_plan] is necessary for the 1%.
An accuracy of 90% or above on PlanBench BlocksWorld [planbench] is necessary for the 1%.
@@ D.5 Adaptation
Test: Achieving fewer than 15 Total Errors on the Wisconsin Card Sorting Test (here) is sufficient for the 1%, subject to memorization and robustness checks. This is related to the ARC-AGI v3 challenge [chollet2019measureintelligence].
@@ E.1.2 Transformation Sequence
Test: A related benchmark is Michelangelo [michelangelo], but it is substantially harder since it uses very long sequences, whereas we use shorter sequences.
@@ E.3.3 Spatial Navigation Memory
System performance on VSI-Bench [vsi_bench] must exceed 80% accuracy.
System performance on MindCube Tiny [mindcube] must exceed 90%.
@@ G.1.6 Figural Fluency
Draw as many unique designs as possible by connecting the dots with exactly four straight lines in 60 seconds [dkefs_design_fluency]:
@@ G.2 Hallucinations
Test: The hallucination rate of the system on SimpleQA [simpleqa] (without tools) must be less than 5%.
@@ H.1 Perception
Note: Image anomaly detection is highly related to the “Odd-Item Out” and the “What’s Missing (WHM)” RIAS-2 subtests [rias2].
@@
What is this? Answer: Airplane [youtube_bb]
What does this depict? Answer: Siberian Husky [imagenet_r]
What does this distorted image depict? Answer: Zebras [imagenet_3dcc]
@@
[flickr30k]
Create a descriptive caption for this. Answer: A ferret rests its head on a black remote control. [flickr30k]
@@
U Intuitivephysics1.mp4 Is this animated scene physically plausible? Answer: No. [intphys2]
U Intuitivephysics2.mp4 Is this animated scene physically plausible? Answer: Yes. [intphys2]
U Running.mov Is this animated scene anomalous? Answer: Yes. [video_anomaly_source]
@@
For image recognition, it is necessary to receive over 85% on ImageNet [imagenet] and 90% on ImageNet-R [imagenet_r].
For video anomaly detection, it is necessary to receive over 85% on IntPhysics 2 [intphys2], a measurement of intuitive physics understanding.